OpenStudy (anonymous):

find the derivative of the function f(x)=e^(2x) ln(x+2)

OpenStudy (anonymous):

product rule: \(\large f=u \cdot v \rightarrow f'=u'v + v'u \) where \(\large u=e^{2x}, \space v=ln(x+2) \) just calculate u' and v' and plug into the formula...